OpenStudy (anonymous):

Guys, please help me. Thanks in advance Find the area of a regular pentagram if one of the sides of the inner pentagon is 3 cm.

OpenStudy (anonymous):

hi @Jper area of pentagon can be calculated by more than one way.. try with the help of the tips in this link.. if it is difficult i will help you http://www.wikihow.com/Find-the-Area-of-a-Pentagon

OpenStudy (anonymous):

For some useful info on the shape's properties:  According to the link, we have the following ratios involved: \[\frac{x}{y}=\frac{y}{x+y}=\frac{y}{z}=\frac{z}{y+z}=\frac{z}{w}\] You can solve for \(y\): \[\begin{align*}\frac{3}{y}&=\frac{y}{3+y}\\\\ 9+3y&=y^2\\\\ y^2-3y-9&=0\\\\ y&=\frac{3\pm\sqrt{45}}{2}\\\\ y&=\frac{3+\sqrt{45}}{2}&\text{ignore the negative root, }y\text{ must be positive} \end{align*}\] From this, you can solve for \(z\) and eventually \(w\): \[\begin{align*} \frac{\dfrac{3+\sqrt{45}}{2}}{3+\dfrac{3+\sqrt{45}}{2}}&=\frac{\dfrac{3+\sqrt{45}}{2}}{z}\\\\ \frac{1}{3+\dfrac{3+\sqrt{45}}{2}}&=\frac{1}{z}\\\\ z&=3+\frac{3+\sqrt{45}}{2}\\\\ z&=\frac{9+\sqrt{45}}{2} \end{align*}\] and in turn, \[\begin{align*} \frac{\dfrac{9+\sqrt{45}}{2}}{\dfrac{3+\sqrt{45}}{2}+\dfrac{9+\sqrt{45}}{2}}&=\frac{\dfrac{9+\sqrt{45}}{2}}{w}\\\\ \frac{1}{6+\sqrt{45}}&=\frac{1}{w}\\\\ w&=6+\sqrt{45}\end{align*}\]